# Corona rigidity

### v3

@@ 4.2. Multiplier and corona algebras
Multiplier algebras provide an abstract generalization of $\mathcal{B}(H)$ and a noncommutative analogue of the Čech–Stone compactification, while corona algebras similarly generalize the Calkin algebra and Čech–Stone remainders of topological spaces (Example 4.4).
@@
Given a $\mathrm{C}^{*}$-algebra $A$, its multiplier algebra ${\mathcal{M}}(A)$ is the unital $\mathrm{C}^{*}$-algebra such that when a unital $\mathrm{C}^{*}$-algebra $B$ contains $A$ as an essential ideal, then the identity map on $A$ extends uniquely to a -homomorphism from $B$ to ${\mathcal{M}}(A)$ ([12, II.7.3.1]). The correspondence with the Čech–Stone compactification is due to the fact that when $A\cong C_{0}(X)$ then ${\mathcal{M}}(C_{0}(X))\cong C(\beta X)$. We refer to [12, II.7.3] and [59, §13] for a rigorous presentation and equivalent definitions of ${\mathcal{M}}(A)$. The following one is illuminating ([59, Corollary 13.2.2]).
@@
For a non-unital $\mathrm{C}^{*}$-algebra $A$, and any nondegenerate [^11] A representation is nondegenerate if $\Phi[A]^{\perp}=\{0\}$. faithful representation $\Phi\colon A\to\mathcal{B}(H)$ there is an isomorphism between ${\mathcal{M}}(A)$ and the idealizer
@@
The multiplier algebra ${\mathcal{M}}(A)$ is the completion of $A$ with respect to the uniformity associated with the strict topology ([59, Lemma 13.1.5]). Moreover, if $A$ is separable, the strict topology is Polish on the unit ball of ${\mathcal{M}}(A)$, giving it the structure of a standard Borel space ([59, Lemma 13.1.8]). If $X$ is a locally compact second countable topological space, this is the topology on $C(\beta X)$ given by uniform convergence on compact subsets of $X$.
@@ 4.2.1. Algebraically trivial automorphisms
We contextualize the discussion in §2 to the category of $\mathrm{C}^{*}$-algebras, aiming to introduce a suitable definition of algebraically trivial isomorphism (Definition 2.3) for corona algebras. A natural notion of triviality for automorphisms of $\mathrm{C}^{*}$-algebras is given by their innerness.
@@
Inner automorphisms of coronas of separable $\mathrm{C}^{*}$-algebras (considered with the Borel structure induced by the strict topology) are topologically trivial. Adopting the notion of inner automorphisms as algebraically trivial ones is however futile in the abelian setting since the only inner automorphism of an abelian $\mathrm{C}^{*}$-algebra is the identity. The definition of algebraically trivial is thus given by generalizing the concept of ‘almost permutation’ (appearing in  [71]) as follows.
@@
The following definition provides a suitable generalization of Definition 4.8 to all $\mathrm{C}^{*}$-algebras.
Definition 4.9 ([170, Definition 2.3]).
@@
In many examples, the notion of algebraically trivial automorphism corresponds to the automorphisms of coronas that are constructed by hand. For abelian $\mathrm{C}^{*}$-algebras, an algebraically trivial isomorphism between $C(\beta X\setminus X)$ and $C(\beta Y\setminus Y)$ is dual to an algebraically trivial homeomorphism as in Definition 4.8 ([170, Proposition 2.7]). Another natural description of algebraically trivial isomorphisms can be made for reduced products of unital separable $\mathrm{C}^{*}$-algebras that do not have central projections (see [170, Proposition 5.3]). In case of the Calkin algebra, algebraic triviality coincides with innerness.
@@
Algebraically trivial isomorphisms are topologically trivial if $A$ and $B$ are separable ([170, Theorem 2.5]). The converse is not known (at least not unconditionally on $A$ and $B$), and it is tightly related to Ulam-stability phenomena. The following is the absolute (see §12) portion of [170, Conjecture 5.1], partially answered in Theorem 5.19 below.
@@ 4.3. The BDF Question
The search for an invariant able to completely classify elements of $\mathcal{B}(H)$ up to essential unitary equivalence goes back to the early investigations on operator algebras by Weyl and von Neumann, where the relation $\sim_{\mathcal{K}(H)}$ naturally came out in the study of the stability of pseudodifferential operators under compact perturbations. In 1909 Weyl proved in [172] that every self-adjoint operator $a$ on a separable Hilbert space can be written as $a=d+c$, where $d$ is a diagonalizable operator and $c\in\mathcal{K}(H)$. Weyl also observed that if $a\sim_{\mathcal{K}(H)}b$ then the essential spectra of $a$ and $b$ are the same. Given an operator $a\in\mathcal{B}(H)$, its essential spectrum $\sigma_{e}(a)$ is the subset of $\mathbb{C}$ consisting of all the elements in the spectrum of $a$ which are either eigenvalues of infinite multiplicity or accumulation points. The set $\sigma_{e}(a)$ is known to be equal to $\sigma(\dot{a})$, the spectrum of $\dot{a}$ in the Calkin algebra ([87, Proposition 2.2.2]).
@@
In Problem 4 of his Ten Problems in Hilbert Space [84], Halmos asked whether every normal operator is the sum of a diagonalizable operator and a compact one, pushing towards an extension of Weyl’s studies. A positive answer to his question was found shortly after by Berg [10] and Sikonia [152], who independently generalized Weyl and von Neumann’s results to normal operators. This collective effort is nowadays summarized by the so called Weyl–von Neumann–Berg–Sikonia Theorem.
@@
This latter characterization also applies to all compact perturbations of normal operators in $\mathcal{B}(H)$. In [20] Brown, Douglas and Fillmore further extended the scope of this classification to all essentially normal operators. These are all the elements $a\in\mathcal{B}(H)$ such that $aa^{*}-a^{*}a\in\mathcal{K}(H)$ or, equivalently, such that $\dot{a}$ is normal in $\mathcal{Q}(H)$.
@@
Brown, Douglas and Fillmore gave a full characterization of unitary equivalence up to compact perturbation on essentially normal operators in [20] by adding the Fredholm index to the invariant.
@@
In  [21] the authors extend their study to $\operatorname{Ext}(X)$ for $X$ a general compact metrizable space, which by the Gelfand–Naimark Duality corresponds to the analysis of $\operatorname{Ext}$ for separable, abelian, unital $\mathrm{C}^{*}$-algebras. They prove that $\operatorname{Ext}(X)$ is always an abelian group, and that it in this setting $\operatorname{Ext}$ is a homotopy invariant, covariant functor, which is moreover a $K$-homology. Their work effectively led to the birth of a brand new subject combining $\mathrm{C}^{*}$-algebras and algebraic topology, known as extension theory (standard references for this subject are [11] and [87]).
@@ 4.4. Sakai’s Question
The methods used to prove this result involve the study of the connected components of the unitary group of $\mathcal{Q}(A)$, which are shown to be directly related to the $K_{0}$-group $K_{0}(A)$, which in turn characterizes $A\otimes\mathcal{K}(H)$ up to isomorphism. We refer to [133, Chapter 1] for further details and proper context on this foundational classification result, and to [173] for an up-to-date overview of the Elliott classification program. In [62, Corollary D] it was proven that $\mathcal{Q}(H)$ is not isomorphic to $\mathcal{Q}(A\otimes\mathcal{K}(H))$ for any Elliott-classifiable $\mathrm{C}^{*}$-algebra $A$, thus giving support to the conjecture that $\mathcal{Q}(H)$ is not isomorphic to $\mathcal{Q}(A)$ for a simple $A$ not isomorphic to the algebra of compact operators. This result was extended in [72, Corollary 5.3] where it was proven that there are $2^{\aleph_{0}}$ non-isomorphic coronas of simple separable $\mathrm{C}^{*}$-algebras.
@@
For an arbitrary $\mathrm{C}^{*}$-algebra $C$ there exists a $\mathrm{C}^{*}$-algebra $A$ such that $\mathcal{Q}(A)\cong C$ (see [74, Proposition 3.2]), but it is not known whether such $A$ can be chosen to be simple. As we said, it is possible to have $\mathcal{Q}(A)\cong\mathbb{C}$ for a simple $A$ ([138, Theorem 1]), and also for an approximately finite (but not simple) $A$ ([81]). Since for a finite-dimensional $\mathrm{C}^{*}$-algebra $F$ and any $A$ we have $F\otimes\mathcal{Q}(A)\cong\mathcal{Q}(F\otimes A)$, the answer to Question 4.21 is positive for a finite-dimensional $C$. Also, for every compact metric space $K$ there is a separable Banach space $X$ such that the ‘Calkin algebra’ $\mathcal{B}(X)/\mathcal{K}(X)$ is isomorphic to $C(K)$ ([115]; the case when $K$ is a singleton was a major open problem until [2]).
@@ 5. Ulam-stability
The present section is of central importance to the subject of corona rigidity, but it has little to do with logic. Readers interested only in logic may want to skip this section and refer to it as necessary. In ([163, §VI.1]) Ulam asked a series of questions about stability of mathematical theorems and solutions to functional equations. In particular, he asked whether ‘approximate’ morphisms in certain categories (such as metric groups or vector spaces) are close to morphisms. The subjects inspired by Ulam’s questions are independently studied and related results have been proved in numerous contexts (see the introduction of [22] for an overview). As we shall see in this chapter, these stability-type questions naturally arise in the study of the rigidity of Borel quotients. In fact, Ulam-stability questions are often tightly connected to whether topologically trivial isomorphisms between quotient structures are algebraically trivial.
@@ 5.2. From Borel to asymptotically additive
In the case when ${\mathcal{I}}=\operatorname{Fin}$, Lemma 5.7 was extracted from Shelah’s proof (§7.1) in [167]. Its proof and the proof of the following (taken from [52, §4]) use the method of stabilizers, isolated by Just in [93].
@@
In the proof, the assumption that all $A_{n}$ are finite is needed in order to use the classical characterization of comeager subsets of the product $\prod_{n}A_{n}$ (see e.g., [59, Theorem 9.9.1]). The analog of Lemma 5.8 is true in some other categories (see §7.4).
@@ 5.3. Boolean algebras
If ${\mathcal{I}}$ is a generalized density ideal (Definition 3.5), then $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ is isomorphic to the reduced product $\prod_{n}(\mathcal{P}(I_{n}),d_{n})/\bigoplus_{\operatorname{Fin}}(\mathcal{P}(I_{n}),d_{n})$ for some finite sets $I_{n}$.
@@
Lemma 5.8, Proposition 5.10, and Proposition 5.11 together imply that, for homomorphisms between quotients over analytic P-ideals (and especially for generalized density ideals), the equivalence of algebraic triviality and topological triviality hinges on Ulam-stability of approximate homomorphisms between metric Boolean algebras of the form $(\mathcal{P}({\mathbb{N}}),d_{\varphi})$ where $d_{\varphi}=\varphi(X\Delta Y)$ for a strictly positive submeasure $\varphi$ on ${\mathbb{N}}$. The failure of Ulam-stability for such class was proved in [49].
@@
An ideal ${\mathcal{I}}$ on ${\mathbb{N}}$ has the Radon–Nikodym property if every topologically trivial Boolean algebra homomorphism $\Phi\colon\mathcal{P}({\mathbb{N}})\to\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ is algebraically trivial (see [50, §1.4 and §1.9]). [^13] The kernel of $\Phi$ is possibly nontrivial.
A submeasure is said to be nonpathological if it is the supremum of measures dominated by it. [^14] This is not the standard definition, according to which no submeasure on a finite set is pathological. Pathological submeasures have been intensively studied in connection to Maharam’s and von Neumann’s problems on characterization of measure algebras, in a work that culminated in [159]. As suggested by the adjective, pathological submeasures are not easy to come by, and for the naturally occurring $F_{\sigma}$ and analytic P-ideals, the submeasure $\varphi$ associated to them (see Theorem 3.6) can usually be chosen to be nonpathological. Such ideals are called nonpathological. In [49] and [51] it was proved that the approximate Boolean algebra homomorphisms and group homomorphisms with respect to nonpathological submeasures are Ulam-stable. Thus Lemma 5.7 and Proposition 5.10 together imply that nonpathological analytic P-ideals have the Radon–Nikodym property. The other parts of the following partial answer to Question 3.7 are proved using similar ideas.
@@
    (((([50, Theorem 1.9.1])))) Nonpathological analytic P-ideals.
@@
    (((([55, Theorem 12.1])))) The ideals $\operatorname{NWD}(\mathbb{Q})$ and $\operatorname{NULL}(\mathbb{Q})$, of subsets of $\mathbb{Q}$ whose closures are nowhere dense and Lebesgue null, respectively and the Weyl ideal ${\mathcal{Z}}_{W}$.
@@
The first four items in Theorem 5.13 apply even when these structures are considered as just groups, with the appropriate notion of Radon–Nikodym property; see [99] and [98] for additional information. Analytic, and even $F_{\sigma}$, ideals without the Radon–Nikodym property exist by [50, Theorem 1.9.5], but the following is open.
@@
A version of this question for analytic P-ideals appears in [50, Question 1.14.3], and its reformulation in terms of Ulam-stability for approximate isomorphisms is [50, Question 1.14.4].

### v4

@@ 4.2. Multiplier and corona algebras
Multiplier algebras provide an abstract generalisation of $\mathcal{B}(H)$ and a noncommutative analogue of the Čech–Stone compactification, while corona algebras similarly generalise the Calkin algebra and Čech–Stone remainders of topological spaces (Example 4.4).
@@
Given a $\mathrm{C}^{*}$-algebra $A$, its multiplier algebra ${\mathcal{M}}(A)$ is the unital $\mathrm{C}^{*}$-algebra such that when a unital $\mathrm{C}^{*}$-algebra $B$ contains $A$ as an essential ideal, then the identity map on $A$ extends uniquely to a -homomorphism from $B$ to ${\mathcal{M}}(A)$ ([12, II.7.3.1]). The correspondence with the Čech–Stone compactification is due to the fact that when $A\cong C_{0}(X)$ then ${\mathcal{M}}(C_{0}(X))\cong C(\beta X)$. We refer to [12, II.7.3] and [61, §13] for a rigorous presentation and equivalent definitions of ${\mathcal{M}}(A)$. The following one is illuminating ([61, Corollary 13.2.2]).
@@
For a non-unital $\mathrm{C}^{*}$-algebra $A$, and any nondegenerate [^11] A representation is nondegenerate if $\Phi[A]^{\perp}=\{0\}$. faithful representation $\Phi\colon A\to\mathcal{B}(H)$ there is an isomorphism between ${\mathcal{M}}(A)$ and the idealiser
@@
The multiplier algebra ${\mathcal{M}}(A)$ is the completion of $A$ with respect to the uniformity associated with the strict topology ([61, Lemma 13.1.5]). Moreover, if $A$ is separable, the strict topology is Polish on the unit ball of ${\mathcal{M}}(A)$, giving it the structure of a standard Borel space ([61, Lemma 13.1.8]). If $X$ is a locally compact second countable topological space, this is the topology on $C(\beta X)$ given by uniform convergence on compact subsets of $X$.
@@ 4.2.1. Algebraically trivial automorphisms
We contextualise the discussion in §2 to the category of $\mathrm{C}^{*}$-algebras, aiming to introduce a suitable definition of algebraically trivial isomorphism (Definition 2.3) for corona algebras. A natural notion of triviality for automorphisms of $\mathrm{C}^{*}$-algebras is given by their innerness.
@@
Inner automorphisms of coronas of separable $\mathrm{C}^{*}$-algebras (considered with the Borel structure induced by the strict topology) are topologically trivial. Adopting the notion of inner automorphisms as algebraically trivial ones is however futile in the abelian setting since the only inner automorphism of an abelian $\mathrm{C}^{*}$-algebra is the identity. The definition of algebraically trivial is thus given by generalising the concept of ‘almost permutation’ (appearing in  [75]) as follows.
@@
The following definition provides a suitable generalisation of Definition 4.8 to all $\mathrm{C}^{*}$-algebras.
Definition 4.9 ([178, Definition 2.3]).
@@
In many examples, the notion of algebraically trivial automorphism corresponds to the automorphisms of coronas that are constructed by hand. For abelian $\mathrm{C}^{*}$-algebras, an algebraically trivial isomorphism between $C(\beta X\setminus X)$ and $C(\beta Y\setminus Y)$ is dual to an algebraically trivial homeomorphism as in Definition 4.8 ([178, Proposition 2.7]). Another natural description of algebraically trivial isomorphisms can be made for reduced products of unital separable $\mathrm{C}^{*}$-algebras that do not have central projections (see [178, Proposition 5.3]). In case of the Calkin algebra, algebraic triviality coincides with innerness.
@@
Algebraically trivial isomorphisms are topologically trivial if $A$ and $B$ are separable ([178, Theorem 2.5]). The converse is not known (at least not unconditionally on $A$ and $B$), and it is tightly related to Ulam-stability phenomena. The following is the absolute (see §12) portion of [178, Conjecture 5.1], partially answered in Theorem 5.19 below.
@@ 4.3. The BDF Question
The search for an invariant able to completely classify elements of $\mathcal{B}(H)$ up to essential unitary equivalence goes back to the early investigations on operator algebras by Weyl and von Neumann, where the relation $\sim_{\mathcal{K}(H)}$ naturally came out in the study of the stability of pseudodifferential operators under compact perturbations. In 1909 Weyl proved in [181] that every self-adjoint operator $a$ on a separable Hilbert space can be written as $a=d+c$, where $d$ is a diagonalisable operator and $c\in\mathcal{K}(H)$. Weyl also observed that if $a\sim_{\mathcal{K}(H)}b$ then the essential spectra of $a$ and $b$ are the same. Given an operator $a\in\mathcal{B}(H)$, its essential spectrum $\sigma_{e}(a)$ is the subset of $\mathbb{C}$ consisting of all the elements in the spectrum of $a$ which are either eigenvalues of infinite multiplicity or accumulation points. The set $\sigma_{e}(a)$ is known to be equal to $\sigma(\dot{a})$, the spectrum of $\dot{a}$ in the Calkin algebra ([91, Proposition 2.2.2]).
@@
In Problem 4 of his Ten Problems in Hilbert Space [88], Halmos asked whether every normal operator is the sum of a diagonalisable operator and a compact one, pushing towards an extension of Weyl’s studies. A positive answer to his question was found shortly after by Berg [10] and Sikonia [159], who independently generalised Weyl and von Neumann’s results to normal operators. This collective effort is nowadays summarised by the so called Weyl–von Neumann–Berg–Sikonia Theorem.
@@
This latter characterisation also applies to all compact perturbations of normal operators in $\mathcal{B}(H)$. In [20] Brown, Douglas and Fillmore further extended the scope of this classification to all essentially normal operators. These are all the elements $a\in\mathcal{B}(H)$ such that $aa^{*}-a^{*}a\in\mathcal{K}(H)$ or, equivalently, such that $\dot{a}$ is normal in $\mathcal{Q}(H)$.
@@
Brown, Douglas and Fillmore gave a full characterisation of unitary equivalence up to compact perturbation on essentially normal operators in [20] by adding the Fredholm index to the invariant.
@@
In  [21] the authors extend their study to $\operatorname{Ext}(X)$ for $X$ a general compact metrisable space, which by the Gelfand–Naimark Duality corresponds to the analysis of $\operatorname{Ext}$ for separable, abelian, unital $\mathrm{C}^{*}$-algebras. They prove that $\operatorname{Ext}(X)$ is always an abelian group, and that it in this setting $\operatorname{Ext}$ is a homotopy invariant, covariant functor, which is moreover a $K$-homology. Their work effectively led to the birth of a brand new subject combining $\mathrm{C}^{*}$-algebras and algebraic topology, known as extension theory (standard references for this subject are [11] and [91]).
@@ 4.4. Sakai’s Question
The methods used to prove this result involve the study of the connected components of the unitary group of $\mathcal{Q}(A)$, which are shown to be directly related to the $K_{0}$-group $K_{0}(A)$, which in turn characterises $A\otimes\mathcal{K}(H)$ up to isomorphism. We refer to [139, Chapter 1] for further details and proper context on this foundational classification result, and to [182] for an up-to-date overview of the Elliott classification program. In [65, Corollary D] it was proven that $\mathcal{Q}(H)$ is not isomorphic to $\mathcal{Q}(A\otimes\mathcal{K}(H))$ for any Elliott-classifiable $\mathrm{C}^{*}$-algebra $A$, thus giving support to the conjecture that $\mathcal{Q}(H)$ is not isomorphic to $\mathcal{Q}(A)$ for a simple $A$ not isomorphic to the algebra of compact operators. This result was extended in [76, Corollary 5.3] where it was proven that there are $2^{\aleph_{0}}$ non-isomorphic coronas of simple separable $\mathrm{C}^{*}$-algebras.
@@
For an arbitrary $\mathrm{C}^{*}$-algebra $C$ there exists a $\mathrm{C}^{*}$-algebra $A$ such that $\mathcal{Q}(A)\cong C$ (see [78, Proposition 3.2]), but it is not known whether such $A$ can be chosen to be simple. As we said, it is possible to have $\mathcal{Q}(A)\cong\mathbb{C}$ for a simple $A$ ([144, Theorem 1]), and also for an approximately finite (but not simple) $A$ ([85]). Since for a finite-dimensional $\mathrm{C}^{*}$-algebra $F$ and any $A$ we have $F\otimes\mathcal{Q}(A)\cong\mathcal{Q}(F\otimes A)$, the answer to Question 4.21 is positive for a finite-dimensional $C$. Also, for every compact metric space $K$ there is a separable Banach space $X$ such that the ‘Calkin algebra’ $\mathcal{B}(X)/\mathcal{K}(X)$ is isomorphic to $C(K)$ ([121]; the case when $K$ is a singleton was a major open problem until [2]).
@@ 5. Ulam-stability
The present section is of central importance to the subject of corona rigidity, but it has little to do with logic. Readers interested only in logic may want to skip this section and refer to it as necessary. In ([170, §VI.1]) Ulam asked a series of questions about stability of mathematical theorems and solutions to functional equations. In particular, he asked whether ‘approximate’ morphisms in certain categories (such as metric groups or vector spaces) are close to morphisms. The subjects inspired by Ulam’s questions are independently studied and related results have been proved in numerous contexts (see the introduction of [22] for an overview). As we shall see in this chapter, these stability-type questions naturally arise in the study of the rigidity of Borel quotients. In fact, Ulam-stability questions are often tightly connected to whether topologically trivial isomorphisms between quotient structures are algebraically trivial.
@@ 5.2. From Borel to asymptotically additive
In the case when ${\mathcal{I}}=\operatorname{Fin}$, Lemma 5.7 was extracted from Shelah’s proof (§7.1) in [174]. Its proof and the proof of the following (taken from [54, §4]) use the method of stabilisers, isolated by Just in [97].
@@
In the proof, the assumption that all $A_{n}$ are finite is needed in order to use the classical characterisation of comeager subsets of the product $\prod_{n}A_{n}$ (see e.g., [61, Theorem 9.9.1]). The analog of Lemma 5.8 is true in some other categories (see §7.4).
@@ 5.3. Boolean algebras
If ${\mathcal{I}}$ is a generalised density ideal (Definition 3.5), then $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ is isomorphic to the reduced product $\prod_{n}(\mathcal{P}(I_{n}),d_{n})/\bigoplus_{\operatorname{Fin}}(\mathcal{P}(I_{n}),d_{n})$ for some finite sets $I_{n}$.
@@
Lemma 5.8, Proposition 5.10, and Proposition 5.11 together imply that, for homomorphisms between quotients over analytic P-ideals (and especially for generalised density ideals), the equivalence of algebraic triviality and topological triviality hinges on Ulam-stability of approximate homomorphisms between metric Boolean algebras of the form $(\mathcal{P}({\mathbb{N}}),d_{\varphi})$ where $d_{\varphi}=\varphi(X\Delta Y)$ for a strictly positive submeasure $\varphi$ on ${\mathbb{N}}$. The failure of Ulam-stability for such class was proved in [51].
@@
An ideal ${\mathcal{I}}$ on ${\mathbb{N}}$ has the Radon–Nikodym property if every topologically trivial Boolean algebra homomorphism $\Phi\colon\mathcal{P}({\mathbb{N}})\to\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ is algebraically trivial (see [52, §1.4 and §1.9]). [^13] The kernel of $\Phi$ is possibly nontrivial.
A submeasure is said to be nonpathological if it is the supremum of measures dominated by it. [^14] This is not the standard definition, according to which no submeasure on a finite set is pathological. Pathological submeasures have been intensively studied in connection to Maharam’s and von Neumann’s problems on characterisation of measure algebras, in a work that culminated in [166]. As suggested by the adjective, pathological submeasures are not easy to come by, and for the naturally occurring $F_{\sigma}$ and analytic P-ideals, the submeasure $\varphi$ associated to them (see Theorem 3.6) can usually be chosen to be nonpathological. Such ideals are called nonpathological. In [51] and [53] it was proved that the approximate Boolean algebra homomorphisms and group homomorphisms with respect to nonpathological submeasures are Ulam-stable. Thus Lemma 5.7 and Proposition 5.10 together imply that nonpathological analytic P-ideals have the Radon–Nikodym property. The other parts of the following partial answer to Question 3.7 are proved using similar ideas.
@@
    (((([52, Theorem 1.9.1])))) Nonpathological analytic P-ideals.
@@
    (((([57, Theorem 12.1])))) The ideals $\operatorname{NWD}(\mathbb{Q})$ and $\operatorname{NULL}(\mathbb{Q})$, of subsets of $\mathbb{Q}$ whose closures are nowhere dense and Lebesgue null, respectively and the Weyl ideal ${\mathcal{Z}}_{W}$.
@@
The first four items in Theorem 5.13 apply even when these structures are considered as just groups, with the appropriate notion of Radon–Nikodym property; see [103] and [102] for additional information. Analytic, and even $F_{\sigma}$, ideals without the Radon–Nikodym property exist by [52, Theorem 1.9.5], but the following is open.
@@
A version of this question for analytic P-ideals appears in [52, Question 1.14.3], and its reformulation in terms of Ulam-stability for approximate isomorphisms is [52, Question 1.14.4].